circular circular electrodes (radius of 3 cm), spaced 1 mm apart, are connected to a 10 V battery. What is the charge on each electrode? 0.25 nC
Answer #1

Its a parallel plate capacitor. Radius of the circular plateis, r 3 cm 0.03 m Distance between the plates is, d-1 mm 10 m Vo

The charge on each electrode is, Q= (2.5×10-11 F)(10 V) 0-2.5×10-10 C= 0.25×10-9 C O-0.25 nC